# Multimodal Outer Arithmetic Block Dual Fusion of Whole Slide Images and Omics Data for Precision Oncology

### v1

@@ Abstract
Developing a central nervous system (CNS) tumor classifier by integrating DNA methylation data with Whole Slide Images (WSI) offers significant potential for enhancing diagnostic precision in neuropathology. Existing approaches typically integrate encoded omic data with histology only once—either at an early or late fusion stage—while reintroducing encoded omic data to create a dual fusion variant remains unexplored. Nevertheless, reintroduction of omic embeddings during early and late fusion enables the capture of complementary information from localized patch-level and holistic slide-level interactions, allowing boosted performance through advanced multimodal integration. To achieve this, we propose a dual fusion framework that integrates omic data at both early and late stages, fully leveraging its diagnostic strength. In the early fusion stage, omic embeddings are projected into a patch-wise latent space, generating omic-WSI embeddings that encapsulate per-patch molecular and morphological insights, effectively incorporating this information into the spatial representation of histology. These embeddings are refined with a multiple instance learning gated attention mechanism to attend to critical patches. In the late fusion stage, we reintroduce the omic data by fusing it with slide-level omic-WSI embeddings using a Multimodal Outer Arithmetic Block (MOAB), which richly intermingles features from both modalities, capturing their global correlations and complementarity. We demonstrate accurate CNS tumor subtyping across 20 fine-grained subtypes and validate our approach on benchmark datasets, achieving improved survival prediction on TCGA-BLCA and competitive performance on TCGA-BRCA compared to state-of-the-art methods. This dual fusion strategy enhances interpretability and classification performance, highlighting its potential for clinical diagnostics.
@@ 1 Introduction
DNA methylation is an epigenetic process where methyl groups are added to specific sites on DNA molecules, typically at cytosine bases followed by guanine (CpG sites) [1]. These modifications affect downstream gene expression levels without changing the DNA sequence itself, resulting in heritable changes in gene activity—an effect termed “epigenetic” [2]. Such a process often leads to gene silencing or reduced gene expression by preventing transcription factors from binding to DNA, thereby inhibiting gene activation. In central nervous system (CNS) tumors, DNA methylation patterns offer valuable insights, enabling the differentiation of tumor subtypes, prediction of clinical outcomes, and guidance on treatment strategies [3]—particularly in cases where diagnostics via digitized Whole Slide Images (WSIs) shows high inter-observer variability [4]. Given the aggressive nature of malignant CNS tumors and their associated poor survival rates, there is a critical need to improve diagnostic precision [5]. One study demonstrated altered subtype classifications in 35% of cases, potentially impacting treatment decisions for 4% of pediatric patients [6]. This demand has driven the development of CNS tumor classifiers based on DNA methylation array data, providing significant advancements in neuropathology [4]. The World Health Organization (WHO) has also responded by incorporating molecular profiling alongside traditional histology into the latest CNS tumor classification guidelines, which defines 40 tumor types and subtypes based on key molecular characteristics [7]. Such findings have further inspired research into automated integrative molecular morphology classification systems using artificial intelligence (AI) algorithms, including machine learning (ML) and deep learning (DL), to improve tumor diagnosis and prognosis [8, 9, 10, 11, 4, 12, 13].
@@
Unimodal Approaches For instance, in the single modality domain, Capper et al. [9] conducted pioneering validation of a methylation-based classification system on 2801 CNS tumor samples using a machine learning (ML) approach [9]. For each CNS tumor subtype included in the classifier, which is referred to as the methylation class, the classifier generated a predicted probability (calibrated score) that summed to 1, with the optimal balance between specificity and sensitivity obtained at 0.84 [14], with tumors with a calibrated score below 0.3 classified as “no match”. Capper’s ML method now serves as an essential aid in the routine diagnostic workup of CNS tumors [4]. Hwang et al. [15] developed an image-based deep learning model using DNA methylation data to predict the origin of cancer of unknown primary (CUP). By employing a vision transformer algorithm and organ-specific DNA methylation images, their approach shows significant potential for enhancing CUP diagnosis and informing treatment strategies. DNA methylation profiling not only aids in precise classification but also supports surgical strategies specific to CNS tumors, improving surgical outcomes and overall patient care [5]. Djirackor et al. [5] utilize ML algorithms to classify brain tumors in real-time by taking the DNA methylation data of a new tumor sample and comparing it to a database of known methylation signatures, assigning a classification based on the closest match. This allows for fast intraoperative decision-making by providing molecular insights during surgery.
@@
Multimodal Approaches. In the multimodal domain, Hoang et al. [12] developed “Deploy”, a deep learning model designed to predict DNA methylation beta values from histopathologic images. Alternatively, Zheng et al. [16] demonstrated that classical machine learning algorithms can link DNA methylation profiles of cancer samples with morphometric features from whole slide images (WSIs), showing improved model performance when genes are grouped into methylation clusters. Interestingly, Sturm et al. [17] explores the use of a multiomic approach — integrating data such as genomics, transcriptomics, and epigenomics — to improve the diagnostic accuracy of pediatric brain tumors, which are often challenging to classify due to overlapping histological features. This method enhances diagnostic precision, allowing for better treatment decisions in pediatric neuro-oncology. However, few studies [12] have combined epigenetic data with WSIs, primarily due to both modalities’ significant data size and complexity, which require extensive preprocessing and advanced fusion techniques to address their heterogeneity. In contrast, multimodal DL approaches combining histology and omic data for improved survival prediction have gained considerable attention in recent years [8, 10, 11, 13, 18, 19, 20, 21, 22, 23, 24, 25]. In particular, several studies [22, 13, 26, 10] have highlighted the value of different fusion stages (early or late), particularly emphasizing early fusion for its ability to create an explainable framework from heterogeneous data. However, Zhang et al. [20] argue that early or late fusion methods can partially overlook modality-specific information, potentially leading to a decline in quantitative or qualitative performance. Accordingly, they propose a prototypical information bottleneck to maintain modality-specific information while simultaneously reduce redundancy. Furthermore, a joint distribution of feature embeddings is used to maintain mutual information between omic and WSI modalities. Despite the promising performance of multimodal approaches in medical diagnostics, significant challenges remain in effectively integrating and analyzing diverse data types, particularly in the context of CNS tumor subtyping. Our critical assessment of existing methods reveals three key gaps:
@@
1) Limitations of WSI: While WSI serves as a primary diagnostic tool for pathologists, accurately identifying fine-grained tumor subtypes based on morphology alone is challenging due to visual feature overlap among CNS subtypes. This similarity risks including irrelevant tumor regions and highlights the need for complementary data sources to achieve more precise subtyping [4]
@@
2) Constraints of DNA methylation classification: DNA methylation classifiers have demonstrated high accuracy in tumor subtyping, but they lack the ability to connect this accuracy to specific regions within WSIs. The main limitation is that these classifiers focus solely on DNA methylation profiles without considering the spatial context provided by WSIs. This constraint limits their capacity to capture how epigenetic patterns contribute to the morphological characteristics observed in specific regions, ultimately reducing the interpretability and comprehensive understanding that could be achieved through integration with WSI data.
@@
3) Scarcity of integrative models for CNS tumor subtyping: This challenge is particularly acute in the context of central nervous system tumors, where the inherent heterogeneity and massive size of both DNA methylation arrays (typically 850k one-dimensional vectors) and WSIs (represented with large, multi-dimensional matrices) have received limited attention. The lack of effective multimodal fusion methods in this domain presents an opportunity to leverage the powerful discriminative capacity of DNA methylation data to improve subtyping and enhance the clinical translation of these advanced imaging and molecular techniques. Addressing these gaps requires innovative approaches to data fusion, potentially exploring both early and late fusion strategies, to fully leverage the complementary strengths of WSI and DNA methylation data in CNS tumor subtyping.
@@ 1.1 Contributions
Answering these challenges, we propose a novel dual fusion approach to improve CNS tumor subtyping by integrating DNA methylation data with WSIs. We design MOAD-Net, a Multimodal Outer Arithmetic Dual Fusion Network that combines two fusion variants: early fusion focused on capturing essential local interactions and late fusion for broader, richer cross-modal correlations, ultimately providing complementary insights and improving the model’s decision-making process. This dual fusion strategy provides a comprehensive, holistic integration of cross-modal data, maximizing the strengths of each fusion type. We summarize the significance of each of our contributions:
  • Importance of early fusion: The early fusion stage enriches interpretability by embedding DNA data directly into each WSI patch, allowing the model to capture detailed, localized interactions. This granularity is essential for identifying unique features within tumor subtypes, especially in cases where subtle variations are critical for accurate subtyping.
  • Significance of late fusion: In contrast, late fusion stage leverages global slide-level associations, capturing high-level, complementary relationships between WSI and omic data.
@@
  • Effectiveness and state-of-the-art results: This combined approach makes MOAD-Net the first method to leverage the BRAIN UK dataset in an imaging-omics pipeline, achieving exceptional results by accurately distinguishing complex brain tumor subtypes. Extensive ablation studies on the TCGA benchmark datasets further validate MOAD-Net’s performance. Our method achieves SOTA survival prediction on the BLCA dataset and competitive results on the BRCA dataset, underscoring its effectiveness across diverse multimodal applications.
@@
We argue that using a late fusion variant alone would capture only global interactions between the WSI level label and epigenetic data, lacking the interpretability needed for accurate subtyping [13, 27]. On the other hand, an early fusion approach alone would focus exclusively on local interactions, neglecting the broader relationships between modalities [27], which we confirmed empirically with extensive experiments. Our proposed dual fusion framework, MOAD-Net, addresses these limitations by combining early and late fusion strategies, creating a new variant that uniquely captures both local and global interactions across modalities and enhances interpretability and overall classification accuracy.
@@
Figure 1: An overview of the proposed MOAD-Net framework. Data engineering and encoding for each modality are performed in the preprocessing block. The early fusion block (top) then receives encoded inputs from both modalities, where omic data is concatenated to form a matrix $z_{i,j}$ which is processed by an MLP encoder that learns a joint mapping, resulting in output $p_{i,j}$ A gated, weighted MIL score is then estimated to indicate patch importance providing heatmap interpretability and producing a WSI feature $W_{i}$. Next, the MOAD-Net fusion block (bottom) reintroduces omic features $o_{i}$ alongside the $W_{i}$ slide-level feature from the early fusion block as input to the MOAB fusion block. This block performs four outer arithmetic operations to create fusion representations, which are further reduced with $f_{\theta}$ before being sent to the final subtyping classifier.
@@
Note that, this paper extends our previous conference paper [28] by implementing a dual fusion architecture, working with three additional imaging-omics datasets, recently proposed SOTA backbones, and shows how heatmaps can be generated, providing interpretability.
@@ 2.1 Datasets
The following sections briefly provide an overview of the datasets used to evaluate MOAD-Net.
2.1.1 Slide Subtyping
UK BRAIN Brain Tumor Dataset[29]: We obtained WSIs and matched DNA methylation data from the UK Brain Archive Information Network (BRAIN UK) [^1] https://www.southampton.ac.uk/brainuk/. The dataset includes (H&E)-stained WSIs from 1,504 patients, covering 20 subtypes, primarily glioblastoma and other glial tumors (left upper corner of Fig. 4). This data exhibits significant class imbalance across its 20 subtypes, with some classes heavily underrepresented, as shown in Fig. 2. To address this, we perform 2-fold cross-validation to ensure all slides in the dataset are covered, and we report the average performance across the two folds for all metrics. For WSIs, we used QuPath [^2] https://qupath.github.io/ for tissue segmentation and tiling, generating 1M patches (4K–10K per slide). For the tissue methylation profile, We processed raw idat files to extract DNA methylation CpG sites using the hg38 Illumina array with EPICv1 annotation, following the workflow in [30]. We calculated m-Values to quantify methylation levels, producing a matrix of 850k CpG sites for 1,504 patients. To reduce features, we applied variance, coefficient of variation (CV), median absolute deviation, and interquartile range, selecting the 8K most variable CpG sites by intersecting these methods. We experimented with clustering 4K CpG sites and Random Forest selection of 10K sites, as suggested in the literature, but 4K reduced performance and 10K showed no improvement, leading us to select 8K CpG sites.
@@ 2.1.1 Slide Subtyping
2.1.2 Survival prediction
TCGA Datasets: To showcase MOAD-Net’s versatility and compare it with other omic-WSI fusion methods [10, 8, 13, 31, 22], we evaluate its performance on The Cancer Genome Atlas (TCGA) datasets for Bladder Urothelial Carcinoma (BLCA) $(n=359)$ and Breast Invasive Carcinoma (BRCA) $(n=869)$, using the survival prediction task outlined by Jaume et al. [13]. We utilized a coupled set of 331 biological pathways derived from 4,999 distinct genes. These genes were grouped into pathways based on their functional interactions and roles in biological processes relevant to the BLCA and BRCA datasets.
2.2 Multimodal Outer Arithmetic Dual Fusion Network (MOAD-Net)
We designed a multimodal fusion framework, shown in Fig. 1, that integrates omic data and WSIs through both early and late fusion stages. Our proposed method, MOAD-Net, is aimed at multimodal brain tumor subtyping, and survival prediction in both lung and breast cancer. In the following subsection, we highlight key components of the MOAD-Net framework.

### v2

@@ Abstract
The integration of DNA methylation data with a Whole Slide Image (WSI) offers significant potential for enhancing the diagnostic precision of central nervous system (CNS) tumor classification in neuropathology. While existing approaches typically integrate encoded omic data with histology at either an early or late fusion stage, the potential of reintroducing omic data through dual fusion remains unexplored. In this paper, we propose the use of omic embeddings during early and late fusion to capture complementary information from local (patch-level) to global (slide-level) interactions, boosting performance through multimodal integration. In the early fusion stage, omic embeddings are projected onto WSI patches in latent-space, which generates embeddings that encapsulate per-patch molecular and morphological insights. This effectively incorporates omic information into the spatial representation of the WSI. These embeddings are then refined with a Multiple Instance Learning gated attention mechanism which attends to diagnostic patches. In the late fusion stage, we reintroduce the omic data by fusing it with slide-level omic-WSI embeddings using a Multimodal Outer Arithmetic Block (MOAB), which richly intermingles features from both modalities, capturing their correlations and complementarity. We demonstrate accurate CNS tumor subtyping across 20 fine-grained subtypes and validate our approach on benchmark datasets, achieving improved survival prediction on TCGA-BLCA and competitive performance on TCGA-BRCA compared to state-of-the-art methods. This dual fusion strategy enhances interpretability and classification performance, highlighting its potential for clinical diagnostics.
@@ 1 Introduction
Epigenetics describes a raft of molecular mechanisms which affect gene expression without changing the DNA sequence itself. DNA methylation is an epigenetic process where methyl groups are added to specific sites on DNA molecules, typically at cytosine bases followed by guanine (CpG sites) [1]. Such a process often leads to gene silencing or reduced gene expression by preventing transcription factors from binding to DNA, thereby inhibiting gene activation, but affects gene expression in many complex ways. The DNA methylation landscape of a cell is dependent on its developmental course and function, and this landscape can be grossly disrupted in the setting of cancer. In central nervous system (CNS) tumors, DNA methylation patterns offer valuable insights, enabling the differentiation of tumor subtypes, prediction of clinical outcomes, and guidance on treatment strategies [2]. Diagnosis solely based on the Whole Slide Image (WSI) can be challenging due to the overlapping appearance of different tumor subtypes, resulting in high inter-observer variability [3].
@@
Given the aggressive nature of malignant CNS tumors and their associated poor survival rates, there is a critical need to improve diagnostic precision [4], expedite diagnostic timeframes and identify targets for future personalized treatments. Importantly, evaluation of digitized WSI is becoming increasingly utilized in these clinical diagnostic pathways. Incorporating DNA methylation profiling for diagnosis, one study in pediatric patients demonstrated altered subtype classifications in 35% of cases, potentially impacting treatment decisions for 4% of pediatric patients [5], while another study in an adult population showed diagnosis was changed in 25%, refined in 4% and confirmed in 25% of cases [6]. This demand has driven the development of CNS tumor classifiers based on DNA methylation array data, providing significant advancements in neuropathology [3]. The World Health Organization (WHO) has also responded by incorporating molecular profiling alongside traditional histology into the latest CNS tumor classification guidelines, which defines 40 tumor types and subtypes based on key molecular characteristics [7]. Such findings have inspired further research into automated integrative molecular morphology classification systems using artificial intelligence (AI) algorithms, including machine learning (ML) and deep learning (DL) approaches, to improve tumor diagnosis and prognosis [8, 9, 10, 11, 3, 12, 13].
@@
Unimodal Approaches. In the single modality domain, Capper et al. [9] pioneered a methylation-based classification system on 2801 CNS tumor samples using a ML approach [9]. For each CNS tumor subtype included in the classifier, also known as the methylation class, the classifier generated a predicted probability (calibrated score) that summed to 1 [14], with tumors with a score below 0.3 classified as “no match”. Capper’s method now serves as an essential aid in the routine diagnostic workup of CNS tumors [3]. Hwang et al. [15] have developed an image-based DL model using DNA methylation data to predict the origin of cancers of unknown primary (CUPs). By employing a vision transformer to organ-specific DNA methylation images, their approach shows significant potential for enhancing CUP diagnosis and informing treatment strategies. DNA methylation profiling not only aids in precise classification but also supports surgical strategies specific to CNS tumors, improving surgical outcomes and overall patient care [4]. Djirackor et al. [4] utilize ML algorithms to classify brain tumors in real-time by taking the DNA methylation data of a new tumor sample and comparing it to a database of known methylation signatures, assigning a classification based on the closest match. This allows for fast intraoperative decision-making by providing molecular insights during surgery.
@@
Multimodal Approaches. In the multimodal domain, Hoang et al. [12] developed “Deploy”, a DL model designed to predict DNA methylation beta values from WSIs. Additionally, Zheng et al. [16] demonstrated that classical ML algorithms can link DNA methylation profiles of cancer samples with morphometric features from WSIs, showing improved model performance when genes are grouped into methylation clusters. Sturm et al. [17] explore the use of a multiomic approach — integrating genomics, transcriptomics, and epigenomics data — to improve the diagnostic accuracy of pediatric brain tumors, which are often challenging to classify due to overlapping histological features. However, few studies [12] have combined epigenetic data with WSIs, primarily due to the significant data size and complexity of both modalities, which require extensive preprocessing and developing advanced fusion techniques to address their heterogeneity.
Multimodal DL approaches combining histology and omic data for improved survival prediction have gained considerable attention in recent years [8, 10, 11, 13, 18, 19, 20, 21, 22, 23, 24, 25]. Several studies [22, 13, 26, 10] have highlighted the value of different fusion stages (early or late), particularly emphasizing early fusion for its ability to create an explainable framework from heterogeneous data. However, Zhang et al. [20] argue that early or late fusion methods can partially overlook modality-specific information, potentially leading to a decline in quantitative or qualitative performance. Accordingly, they propose a prototypical information bottleneck framework to maintain modality-specific information while simultaneously reducing redundancy. Furthermore, a joint distribution of feature embeddings is used to calculate the mutual information between omic and WSI modalities.
Despite the promising performance of multimodal approaches in medical diagnostics, significant challenges remain in effectively integrating and analyzing diverse data types, particularly in the context of CNS tumor subtyping. Our assessment of existing methods reveals three key gaps:
@@
1) Limitations of WSI-only diagnosis: While WSI serves as a primary diagnostic tool for pathologists, accurately identifying fine-grained tumor subtypes based on morphology alone is challenging due to visual feature overlap among CNS subtypes. This similarity increases the risk of including irrelevant tumor regions and highlights the need for complementary data sources to achieve more precise subtyping [3].
@@
2) Shortcomings of DNA methylation classification: DNA methylation classifiers have demonstrated high accuracy in tumor subtyping, but they lack the ability to connect this accuracy to specific regions within WSIs. The main limitation is that these classifiers focus solely on DNA methylation profiles without considering the spatial context provided by WSIs. This limits their capacity to capture how epigenetic patterns contribute to the morphological characteristics observed in specific regions, ultimately reducing the interpretability and comprehensive understanding that could be achieved through integration with WSI data.
@@
3) Scarcity of integrative models for CNS tumor subtyping: This challenge is particularly acute in the context of central nervous system tumors, where the inherent heterogeneity and large size of both DNA methylation arrays (typically 850k one-dimensional vectors) and WSIs (represented with multi-dimensional matrices up to 150k x 150k pixels) have received limited attention. The lack of effective multimodal fusion methods in this domain presents an opportunity to leverage the powerful discriminative capacity of DNA methylation data to improve subtyping and enhance the clinical translation of these advanced imaging and molecular techniques.
@@ 1.1 Contributions
Addressing these challenges, we propose a novel dual fusion approach to improve CNS tumor subtyping by integrating DNA methylation data with WSIs. We design MOAD-FNet, a Multimodal Outer Arithmetic Dual Fusion Network that combines two fusion variants: early fusion focused on capturing essential local interactions and late fusion for broader, richer cross-modal global context, ultimately providing complementary insights and improving the model’s decision-making process. This dual fusion strategy provides a comprehensive, holistic integration of cross-modal data, maximizing the strengths of each fusion type. Our main contributions are as follows:
  • We introduce a novel dual fusion network that seamlessly incorporates both early and late fusion approaches, enabling detailed integration of molecular and imaging data at patch and slide levels.
  • Our approach develops a unique Multimodal Outer Arithmetic Block (MOAB) fusion strategy that enhances cross-modal feature interaction and improves the model’s ability to capture complex tumor subtype features.
@@
  • Our method, MOAD-FNet, is the first imaging-omics pipeline to leverage the NHNN BRAIN UK dataset, demonstrating exceptional performance in brain tumor subtyping. Extensive ablation studies on TCGA datasets validate its robustness, achieving state-of-the-art survival prediction on the BLCA dataset and on par results on the BRCA dataset, showcasing its versatility across multimodal oncology tasks.
@@
We argue that using a late fusion variant alone would capture only global interactions between the WSI level label and epigenetic data, lacking the interpretability needed for accurate subtyping [13, 27]. On the other hand, an early fusion approach alone would focus exclusively on local interactions, neglecting the broader relationships between modalities [27], which we confirm with extensive experimentation. Our proposed dual fusion framework, MOAD-FNet, addresses these limitations by combining early and late fusion strategies, creating a new approach that uniquely captures both local and global interactions across modalities and enhances interpretability and overall classification accuracy.
@@
Figure 1: Overview of the proposed MOAD-FNet framework. Data engineering and encoding for each modality are performed in the preprocessing block. The early fusion block (top) receives encoded inputs from both modalities, where omic data is concatenated to form a matrix $\mathbf{z}_{i,j}$ which is processed by an MLP encoder that learns a joint mapping, resulting in output $\mathbf{p}_{i,j}$. A gated attention via multiple instance learning (MIL) scores patch importance providing heatmap interpretability and producing a WSI feature $\mathbf{v}_{i}$. Next, the MOAD-FNet fusion block (bottom) reintroduces omic features $\mathbf{o}_{i}$ alongside the $\mathbf{v}_{i}$ feature representation from the early fusion block as input to the MOAB fusion block. This block performs four outer arithmetic operations to create fusion representations, which are further reduced with $f_{\theta}$ before being sent to the final subtyping classifier.
@@
Note that, this paper extends our previous conference paper [28] by implementing a dual fusion architecture, working with three additional imaging-omics datasets, recently proposed state-of-the-art (SOTA) backbones, and shows how heatmaps can be generated, providing interpretability.
@@ 2.1 Datasets
The following sections briefly provide an overview of the datasets used to evaluate MOAD-FNet.
@@
2.1.1 Slide Subtyping - NHNN BRAIN UK Brain Tumor Dataset
We obtained WSIs and matched DNA methylation data from the NHNN, University College London Hospital (UCLH) through the UK Brain Archive Information Network [29] [^1] https://www.southampton.ac.uk/brainuk. The dataset includes (H&E)-stained WSIs from 1,504 patients, covering 20 DNA-methylation-based subtypes of high- and low-grade glial tumors. The data exhibits significant class imbalance across its 20 subtypes, as shown in Fig. 2. To ensure a balanced evaluation, we perform 2-fold cross-validation by splitting the dataset into two equal parts (50%-50%). In each fold, one part is used for training and the other for testing, ensuring all slides are evaluated while maintaining a strict separation between the two sets. We report the average performance across both folds for all metrics. For WSIs, we used QuPath [30] for tissue segmentation and tiling, generating 1M patches (4K–10K per slide). For the tissue methylation profile, we processed raw idat files to extract DNA methylation CpG sites using the hg38 Illumina array with EPICv1 annotation, following the workflow in [31]. We calculated m-Values to quantify methylation levels, producing a matrix of 850k CpG sites for 1,504 patients. To reduce features, we applied variance, coefficient of variation (CV), median absolute deviation, and inter-quartile range, selecting the 8K most variable CpG sites by intersecting these methods. We experimented with clustering 4K CpG sites and Random Forest selection of 10K sites, as suggested in the literature[9, 31, 32], but 4K reduced performance and 10K showed no improvement, leading us to select 8K CpG sites.
2.1.2 Survival prediction - TCGA Datasets
To showcase MOAD-FNet’s versatility and compare it with other omic-WSI fusion methods [10, 8, 13, 33, 22], we evaluate its performance on The Cancer Genome Atlas (TCGA) datasets for Bladder Urothelial Carcinoma (BLCA) $(n=359)$ and Breast Invasive Carcinoma (BRCA) $(n=869)$, using the survival prediction task outlined by Jaume et al. [13]. We utilized a coupled set of 331 biological pathways derived from 4,999 distinct genes. These genes were grouped into pathways based on their functional interactions and roles in biological processes relevant to the BLCA and BRCA datasets.
2.2 Multimodal Outer Arithmetic Dual Fusion Network
We designed a multimodal fusion framework, shown in Fig. 1, that integrates omic data and a WSI through combined early and late fusion stages. Our proposed method, MOAD-FNet, is aimed at multimodal brain tumor subtyping, and survival prediction in both lung and breast cancer. In the following subsections, we highlight key components of the MOAD-FNet framework.